OpenStudy (anonymous):

Evaluate the following integral: e^(3x) (e^(x) + 1) dx between ln5 and ln2.

OpenStudy (anonymous):

\[\int\limits_{\ln2}^{\ln5}e^{3x}(e^x+1)dx\]\[=\int\limits_{\ln2}^{\ln5}(e^{4x}+e^{3x})dx\]\[=\frac{e^{4x}}{4}+ \frac{e^{3x}}{3}\]now u evaluate that from ln2 to ln5